# Improving the Cosmological Constraints by Inferring the Formation Channel of Extreme-mass-ratio Inspirals

### v1

@@ 2.3 Statistical framework for constraining the formation channel of EMRIs
To statistically test our null hypothesis, we use the same statistical framework presented in Zhu & Chen (2024), which was developed upon the work of Bartos et al. (2017) and modified to test the spatial correlation between MBHB mergers and AGNs. Given that the spatial number density of AGNs is $\rho_{\rm agn}$, and the spatial localization error volume of the $i$th detected EMRI is $\Delta V_{i}$, in the null hypothesis the number $N_{{\rm agn},i}$ of AGNs within the error volume $\Delta V_{i}$ will follow a Poisson distribution with an expectation of $\rho_{\rm agn}\Delta V_{i}$, i.e.,
@@
(Zhu & Chen, 2024), where $f_{{\rm compl},i}^{\rm agn}$ represents the completeness fraction of the AGN catalog relative to the $i$th detected EMRI. For the estimation of $f_{{\rm compl},i}^{\rm agn}$, we follow exactly the approach of Zhu & Chen (2024), by considering the correlation between the luminosities of AGNs and the masses of their central MBHs. For a GW catalog composed of $N$ EMRIs, the total likelihood can be expressed as
@@
Finally, to statistically constrain the fraction $f_{\rm agn}$ in Equation (17), we follow Bartos et al. (2017) to introduce a likelihood ratio
@@
In the null hypothesis, in which the value of $f_{\rm agn}$ equals to the fraction of AGNs among all galaxies in the Universe, the likelihood ratio $\lambda$ is governed by the random fluctuation of the spatial number density of AGNs, and will follow a “background distribution” $P_{\rm bg}(\lambda)$. A very important property is that the median of $\lambda$ derived from the detected EMRIs will deviate more from $P_{\rm bg}(\lambda)$ when the fraction of the EMRIs originate in AGNs is larger. In the analysis, $P_{\rm bg}(\lambda)$ can be determined by Monte-Carlo simulations, and the deviation of median $\lambda$ relative to $P_{\rm bg}(\lambda)$ can be quantified by an $p$-value. If this $p$-value becomes less than $0.00135$, we can claim that we can reject the null hypothesis with a significance of $3\sigma$, i.e., confirm the spatial correlation between the EMRIs and AGNs.
@@ 3.1 Population models of EMRIs
The EMRI rate is the primary feature of the EMRI population, which is mainly determined by four factors (Babak et al., 2017): (i) the properties of the MBH population, including the mass function and spin distribution of MBHs; (ii) the $M-\sigma$ relation, including the properties of the stellar didtribution surrounding the MBH, the presence and regrowth time of cusp erosion following MBHB formation and evolution, and the EMRI duty cycle; (iii) the ratio of plunges to EMRIs; (iv) the characteristic masses of inspiralling stellar-mass compact objects. All of these factors are uncertain in astrophysics, and different conditions produce different EMRI population models. In this work, we adopt 11 EMRI population models proposed in Babak et al. (2017), i.e., from M1 to M12 (except the M11 model, which predicts a very low EMRI rate so that LISA and TianQin detecting almost no EMRIs, e.g., Babak et al., 2017; Fan et al., 2020). We set the eccentricities of EMRIs at plunge to be uniformly distributed in the range from 00 to $0.2$. The main differences between the various models are summarized in the Table I of Babak et al. (2017).
@@ 3.2 Mock observations
To estimate the abilities of detectors in constraining the cosmic expansion history and the formation channel of EMRIs, we need to simulate (i) the detected EMRI catalogs, and (ii) the redshift probability distribution of the candidate host galaxies as well as (iii) the number of candidate host AGNs corresponding to each detected EMRI.
To generate the detected EMRI catalogs, we follow the approach of Babak et al. (2017) and Fan et al. (2020). First, we adopt a galaxy catalog from the MultiDark cosmological simulations (Klypin et al., 2016; Croton et al., 2016) to select the host galaxies of the EMRIs, with the mass of the MBH at the host galaxy center concordant with the mass of the MBH of the corresponding EMRI. Second, we generate the GW signals of EMRIs using the analytic kludge Kerr (AKK) waveform model (Barack & Cutler, 2004; Babak et al., 2017). The reason we adopt the AKK waveform model here is to facilitate comparison with the previous works (MacLeod & Hogan, 2008; Laghi et al., 2021), although several new waveform models (Chua et al., 2017, 2021; Hughes et al., 2021; Isoyama et al., 2022) have been proposed recently. Third, we calculate the detector responses to the EMRI GW signals, and consider a total of three detector configurations, namely
@@
  • TianQin+LISA (TQ+LISA): a network composed of TianQin and LISA, the LISA configuration according to Amaro-Seoane et al. (2017). In this work, we assume that LISA and TianQin have a five year mission overlap and only consider EMRIs that are individually detectable by both TianQin and LISA.
@@
Figure 1: Panels (a), (b), (c), and (d) show, respectively, the distributions of the total $D_{L}$ estimation errors, the sky localization errors, the numbers of the candidate host galaxies, and the expected numbers of the candidate host AGNs, for detectable EMRIs under the M1 population model. The colors indicate TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). And in Panels (c) and (d), the solid and dashed lines, respectively, represent the conditions in which the cosmological parameters are fixed and freely varying within a prior range, when transforming the ranges of luminosity distances of EMRIs into the ranges of redshifts.
Here, we take the M1 model as the fiducial model as in Laghi et al. (2021). For example, the distribution of the spatial localization errors (including the $D_{L}$ estimation and sky localization errors) of various detector configurations are shown in Panels (a) and (b) of Figure 1. We find that the medians of the estimation errors for luminosity distance are about $7\%$, $6\%$, and $3\%$ for TianQin, TianQin I+II, and TianQin+LISA, respectively, after accounting for additional systematic errors due to the weak lensing effects (Hirata et al., 2010; Cusin & Tamanini, 2021) and peculiar velocities of the host galaxies (Kocsis et al., 2006). The corresponding medians errors for sky localization are about $2.6~{}\!{\rm deg}^{2}$, $0.8~{}\!{\rm deg}^{2}$, and $0.02~{}\!{\rm deg}^{2}$, respectively.
For the cosmological inferences, in order to simulate the redshift probability distributions of the detected EMRIs, we select their candidate host galaxies from the MultiDark galaxy catalog for which the selection effects of surveys have been considered. Here we set the limiting magnitude of the MultiDark galaxy catalog to $m_{\rm limit}^{*}=+26~{}\!{\rm mag}$ according to the survey depth of the Chinese Space Station Telescope (CSST; Gong et al., 2019). To select the candidate host galaxies, we adopt a $2\sigma$ confidence interval (CI) of the spatial localization error volumes, and convert the range of luminosity distance, $[D_{L;\min},D_{L;\max}]$, into a range of redshift, $[z_{\min},z_{\max}]$, using the relationship
@@
The distributions of the numbers of the candidate host galaxies are shown in Panel (c) of Figure 1. We can see that the typical number of candidate host galaxies for an EMRI detected by TianQin or TianQin I+II is about $10^{5}-10^{6}$ after accounting for the uncertainties in the cosmological parameters, and that the TianQin+LISA network can reduce the typical number to about $10^{3}-10^{4}$. In addition, to make the mock galaxy catalog more realistic, we assume that all galaxies with $z>0.6$ in the MultiDark galaxy catalog have photometric redshift (photo-$z$) errors which vary with the true redshifts $z_{\rm true}$ of the galaxies as
@@
represents the observational limiting absolute magnitude for the candidate host AGN of the $i$th EMRI at a distance of $D_{L,i}$, and $BC$ is a factor for bolometric correction (Duras et al., 2020). In this work, we set the limit magnitude of the AGN survey to $m_{\rm limit}^{*{\rm agn}}=+24.4~{}\!{\rm mag}$ according to the depth in the shallowest band of the CSST survey as well as taking into account the need for the spectroscopic identifications of AGNs (Gong et al., 2019). Taking the M1 model as an example, the distributions of the numbers of candidate host AGNs of the EMRIs for various detector configurations are shown in Panel (d) of Figure 1. For the same EMRIs, the numbers of candidate host AGNs are about $1\%$ of the numbers of the normal galaxies which are candidate hosts. This fraction is consistent with the AGN fraction among all galaxies.
4 Constraints on the cosmological parameters
@@ 4.1 Constraints on $H_{0}$ and $\Omega_{M}$
We first fit the $D_{L}-z$ relation, using the luminosity distances from EMRI observations and the redshifts extracted from galaxy surveys. The result is shown in Figure 2, and the corresponding constraints on $H_{0}$ and $\Omega_{M}$ are shown in Figure 3. Note that n this section we fix the parameters of the dark energy EoS to $w_{0}=-1$ and $w_{a}=0$.
Figure 2: Typical example of fitting the $D_{L}-z$ relation derived from the EMRI dark sirens which are detectable by the TianQin+LISA network. Here we use the fiducial EMRI population model. The horizontal colored serial squares represent the probability distribution of redshift (normalized by the maximum) derived from the observed candidate host galaxies, and the color hues from white to red show the probability of redshift. The luminosity distances shown here represent the mean values, and their errors ($1\sigma$ CI) are indicated by the black error bars. The solid green line represents the true cosmological model adopted by our simulations. The dotted-dashed red line represents the best-fit cosmological model, while the two shaded regions representing the $1\sigma$ and $2\sigma$ CIs.
Figure 3: Constraining errors of $h$ and $\Omega_{M}$ for TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). The horizontal axis represents the different EMRI population models. The error bars correspond to a $1\sigma$ CI. The gray shade represents the $1\sigma$ equivalent error of the prior, and the green shade represents a error scale of $2\%$ ($10\%$) for $h$ ($\Omega_{M}$).
For the Hubble-Lemaître constant $H_{0}$, under the fiducial population model, i.e., the M1 model, the expected error is about $8.1\%$ for TianQin, and TianQin I+II can reduce this error to about $4.4\%$. Under the other population models of from M2 to M12, the error of $H_{0}$ varies considerably from model to model. The variations are mainly caused by the fact that the detection rates of EMRIs predicted by various models vary considerably (see Babak et al., 2017; Fan et al., 2020). In general, TianQin is able to achieve a precision in the range of about $3.4\%-9.6\%$, while that for TianQin I+II is expected to be in the range of about $1.5\%-7.2\%$.
@@
If TianQin and LISA could form a network, i.e., TianQin+LISA, the constraints on $H_{0}$ and $\Omega_{M}$ are significantly improved in all population models, as is shown by the red error bars in Figure 3. For example, in the M1 model, the errors of $H_{0}$ and $\Omega_{M}$ can shrink to about $1.9\%$ and $20\%$, respectively. Under the other models, the precision of $H_{0}$ can be improved to about $0.4\%-7.1\%$, and the error of $\Omega_{M}$ would reduce to the range of about $13\%-43\%$. For clarity, we summarize the expected precisions on $H_{0}$ and $\Omega_{M}$ for various cases in Table I.

### v2

@@ 2.3 Statistical framework for constraining the formation channel of EMRIs
To statistically test our null hypothesis, we use the same statistical framework presented in Zhu & Chen (2024), which was developed upon the work of (Bartos et al., 2017, also see earlier Braun et al. (2008)) and modified to test the spatial correlation between MBHB mergers and AGNs. Given that the spatial number density of AGNs is $\rho_{\rm agn}$, and the spatial localization error volume of the $i$th detected EMRI is $\Delta V_{i}$, in the null hypothesis the number $N_{{\rm agn},i}$ of AGNs within the error volume $\Delta V_{i}$ will follow a Poisson distribution with an expectation of $\rho_{\rm agn}\Delta V_{i}$, i.e.,
@@
(Zhu & Chen, 2024), where $f_{{\rm compl},i}^{\rm agn}$ represents the completeness fraction of the AGN catalog when the $i$th detected EMRI is evaluated. For the estimation of $f_{{\rm compl},i}^{\rm agn}$, we follow exactly the approach of Zhu & Chen (2024), by considering the correlation between the luminosities of AGNs and the masses of their central MBHs. For a GW catalog composed of $N$ EMRIs, the total likelihood can be expressed as
@@
Several factors render it exceedingly challenging to accurately infer $f_{\rm agn}$ directly from Equation (17), including (i) the degeneracy between $f_{\rm agn}$ and $f_{{\rm compl},i}^{\rm agn}$, (ii) the sensitivity of the likelihood to the accurate determination of $\rho_{\rm agn}$, and (iii) the data $\{N_{{\rm agn},i}\}$ of the likelihood is not directly related to $f_{\rm agn}$. Regarding (i) and (ii), accurate estimations of both $f_{{\rm compl},i}^{\rm agn}$ and $\rho_{\rm agn}$ are impeded by selection effects. For (iii), Because the number $N_{{\rm agn}}$ of candidate host AGNs is proportional to the spatial localization error volume of the EMRI, not by correlation with $f_{\rm agn}$. An EMRI originated in an AGN merely implies an anticipated increase of one in the count of candidate host AGNs for this EMRI, compared to a random scenario.
To statistically constrain the fraction $f_{\rm agn}$ in Equation (17), we follow Braun et al. (2008) and Bartos et al. (2017) to introduce a likelihood ratio
@@
In the null hypothesis, in which the value of $f_{\rm agn}$ equals to the fraction of AGNs among all galaxies in the universe, the likelihood ratio $\lambda$ is governed by the random fluctuation of the spatial number density of AGNs, and will follow a “background distribution” $P_{\rm bg}(\lambda)$. Conversely, in the alternative hypothesis that $f_{\rm agn}$ significantly exceeds the fraction of AGNs among all galaxies, $\lambda$ becomes governed by the spatial correlation between EMRIs and AGNs. This is because, under this condition, the likelihoods $\mathcal{L}(f_{\rm agn})$ and $\mathcal{L}(0)$ in Equation (18) exhibit significant discrepancies. For illustrative purposes, the distributions of the likelihood ratio $\lambda$ in the null and alternative hypotheses, as well as their variation with $f_{\rm agn}$, are presented in Appendix A. A very important property is that the $\lambda$ derived from real observation of EMRIs will deviate more from $P_{\rm bg}(\lambda)$ when the fraction of the EMRIs originate in AGNs is larger. In the analysis, $P_{\rm bg}(\lambda)$ can be determined by Monte-Carlo simulations, and the deviation of the median $\lambda$ relative to $P_{\rm bg}(\lambda)$ can be quantified by an $p$-value. If this $p$-value becomes less than $0.00135$, we can claim that we can reject the null hypothesis with a significance of $3\sigma$, i.e., confirm the spatial correlation between the EMRIs and AGNs.
Finally, we estimate the uncertainty in the observational constraint on $f_{\rm agn}$. We denote the likelihood ratio derived from observational data (the numbers of candidate host AGNs of the detected EMRIs) as $\lambda_{\rm s}(f_{\rm agn})$, where the subscript “s” stands for “signal”. Notice that its value varies with the underlying AGN fraction $f_{\rm agn}$. To check whether $\lambda_{\rm s}(f_{\rm agn})$ is consistent with or can be rejected by the background distribution of $\lambda$, we introduce an integrated probability
$CDF(\lambda_{\rm s},f_{\rm agn})=\frac{1}{\mathcal{N}}\int_{\lambda_{\rm s}(f_{\rm agn})}^{\infty}\!\!P(\lambda|f_{\rm agn})\mathrm{d}\lambda,$ (19)
where $P(\lambda|f_{\rm agn})$ is the conditional PDF of the likelihood ratio $\lambda$ given $f_{\rm agn}$, and $\mathcal{N}$ is a normalization factor. In our simulations, each $P(\lambda|f_{\rm agn})$ is derived from about $10^{4}$ independent Monte-Carlo simulations.
The meanings of $\lambda_{\rm s}(f_{\rm agn})$ and $P(\lambda|f_{\rm agn})$ are illustrated in Figure 1. It shows that the conditional PDF $P(\lambda|0.1)$ (black dotted curve) is clearly inconsistent with the observed value of $\lambda_{\rm s}$, because $CDF(\lambda_{\rm s},0.1)$ is small. As $f_{\rm agn}$ increases, the PDF $P(\lambda|f_{\rm agn})$ shifts towards larger values of $\lambda$, and becomes more consistent with $\lambda_{s}$. Therefore, we estimate the error of $f_{\rm agn}$ through the following steps. (i) We increase $f_{\rm agn}$ until we find a value which satisfies $CDF(\lambda_{\rm s},f^{\prime}_{\rm agn})=0.15865(0.02275)$. Any $f_{\rm agn}<f^{\prime}_{\rm agn}$ should lie outside the $1\sigma(2\sigma)$ confidence interval (CI) of the probability distribution of $f_{\rm agn}$. (ii) We keep increasing $f_{\rm agn}$ until we find $CDF(\lambda_{\rm s},f^{\prime\prime}_{\rm agn})=0.5$. This $f^{\prime\prime}_{\rm agn}$ represents the median value of $f_{\rm agn}$. (iii) We further increase $f_{\rm agn}$ until $CDF(\lambda_{\rm s},f^{\prime\prime\prime}_{\rm agn})>1-0.15865(0.02275)$. Such $f^{\prime\prime\prime}_{\rm agn}$ determines the $1\sigma(2\sigma)$ upper bound of $f_{\rm agn}$.
Figure 1: An example comparing the signal likelihood ratio $\lambda_{\rm s}(f_{\rm agn})$ (vertical green dashed line) with the conditional PDFs of $\lambda$ for four different values of $f_{\rm agn}$. In this example, $f^{0}_{\rm agn}=0.1$, $f^{\prime}_{\rm agn}=0.55$, $f^{\prime\prime}_{\rm agn}=0.7$, and $f^{\prime\prime\prime}_{\rm agn}=0.85$.
@@ 3.1 Population models of EMRIs
The EMRI rate is a primary feature of the EMRI population, and it is mainly determined by four factors (Babak et al., 2017): (i) the properties of the MBH population, including the mass function and spin distribution of MBHs; (ii) the $M-\sigma$ relation; (iii) the fraction of plunges among EMRIs; (iv) the characteristic masses of inspiralling stellar-mass compact objects. All of these factors are uncertain in astrophysics, and different conditions produce different EMRI population models. In this work, we adopt 11 EMRI population models proposed in Babak et al. (2017), i.e., from M1 to M12 (except the M11 model, which predicts a very low EMRI rate so that LISA and TianQin detecting almost no EMRIs, e.g., Babak et al., 2017; Fan et al., 2020). We set the eccentricities of EMRIs at plunge to be uniformly distributed in the range from 00 to $0.2$. The main differences between the various models are summarized in Table I of Babak et al. (2017). When simulating the EMRIs originating in AGNs, we adjust the corresponding source parameters based on the population models presented in Pan et al. (2021), for instance, by updating the redshift distribution of EMRIs and setting all eccentricities at plunge to zero.
@@ 3.2 Mock observations
To estimate the abilities of detectors in constraining the cosmic expansion history and the formation channel of EMRIs, we need to simulate (i) the detected EMRI catalog, (ii) the redshift probability distribution of the candidate host galaxies, and (iii) the number of candidate host AGNs corresponding to each detected EMRI.
To generate a catalog of detected EMRIs, we follow the approach of Babak et al. (2017) and Fan et al. (2020). First, we adopt a galaxy catalog from the MultiDark cosmological simulations (Klypin et al., 2016; Croton et al., 2016) to select the host galaxies of the EMRIs. In the selection process, the mass of the MBH at the center of the selected host galaxy matches the mass of the MBH of the corresponding EMRI. Second, we generate the GW signals of EMRIs using the analytic kludge Kerr (AKK) waveform model (Barack & Cutler, 2004; Babak et al., 2017). The reason for us to adopt this model is to facilitate comparison with the previous works (MacLeod & Hogan, 2008; Laghi et al., 2021), although several new waveform models (Chua et al., 2017, 2021; Hughes et al., 2021; Isoyama et al., 2022) have been proposed recently. Third, we calculate the detector responses to the EMRI GW signals.
We consider three detector configurations, namely
@@
  • TianQin+LISA (TQ+LISA): a network composed of TianQin and LISA, with the LISA configuration according to Amaro-Seoane et al. (2017). In this work, we assume that LISA and TianQin have a five year mission overlap and only consider EMRIs that are individually detectable by both TianQin and LISA. Here we emphasize that because both LISA (Colpi et al., 2024) and TianQin (Mei et al., 2021) are planned to be launched around 2035, it is possible that there will be a window of five years of overlap.
@@
Figure 2: Panels (a), (b), (c), and (d) show, respectively, the distributions of the total $D_{L}$ estimation errors, the sky localization errors, the numbers of the candidate host galaxies, and the expected numbers of the candidate host AGNs. All results are derived based on the M1 population model. The colors indicate TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). And in Panels (c) and (d), the solid and dashed lines, respectively, represent the conditions in which the cosmological parameters are fixed and freely varying within a prior range, when transforming the range of luminosity distance into the range of redshift.
For example, when using the M1 model (one can compare our results with those in Laghi et al., 2021), the distribution of the spatial localization errors (including the $D_{L}$ estimation and sky localization errors) of various detector configurations are shown in Panels (a) and (b) of Figure 2. We find that the medians of the estimation errors for luminosity distance are about $7\%$, $6\%$, and $3\%$ for TianQin, TianQin I+II, and TianQin+LISA, respectively, after accounting for additional systematic errors due to the weak lensing effects (Hirata et al., 2010; Cusin & Tamanini, 2021) and peculiar velocities of the host galaxies (Kocsis et al., 2006). The corresponding medians errors for sky localization are about $2.6~{}\!{\rm deg}^{2}$, $0.8~{}\!{\rm deg}^{2}$, and $0.02~{}\!{\rm deg}^{2}$, respectively.
For the cosmological inferences, in order to simulate the redshift probability distributions of the detected EMRIs, we select their candidate host galaxies from the MultiDark galaxy catalog, and we take into account the selection effects of surveys. Here we set the limiting magnitude of the MultiDark galaxy catalog to $m_{\rm limit}^{*}=+26~{}\!{\rm mag}$ according to the survey depth of the Chinese Space Station Telescope (CSST; Gong et al., 2019). To select the candidate host galaxies, we adopt a $2\sigma$ CI of the spatial localization error volumes, and convert the range of luminosity distance, $[D_{L;\min},D_{L;\max}]$, into a range of redshift, $[z_{\min},z_{\max}]$, using the relationship
@@
The distribution of the number of the candidate host galaxies is shown in Panel (c) of Figure 2. We can see that the typical number of candidate host galaxies for an EMRI detected by TianQin or TianQin I+II is about $10^{5}-10^{6}$ after accounting for the uncertainties in the cosmological parameters, and that the TianQin+LISA network can reduce the typical number to about $10^{3}-10^{4}$. In addition, to make the mock galaxy catalog more realistic, we assume that all galaxies with $z>0.6$ in the MultiDark galaxy catalog have photometric redshift (photo-$z$) errors which vary with the true redshifts $z_{\rm true}$ of the galaxies as
@@
represents the observational limiting absolute magnitude for the candidate host AGN of the $i$th EMRI at a distance of $D_{L,i}$, and $BC$ is a factor for bolometric correction (Duras et al., 2020). In this work, we set the limiting magnitude of the AGN survey to $m_{\rm limit}^{*{\rm agn}}=+24.4~{}\!{\rm mag}$ according to the depth in the shallowest band of the CSST survey as well as taking into account the need for the spectroscopic identifications of AGNs (Gong et al., 2019). Taking the M1 model as an example, the distributions of the numbers of candidate host AGNs of the EMRIs for various detector configurations are shown in Panel (d) of Figure 2. We can see that for the same EMRI, the number of candidate host AGNs is about $1\%$ of the number of the normal galaxies which are candidate hosts. This fraction is consistent with the AGN fraction among all galaxies.
4 Constraints on cosmological parameters
@@ 4.1 Constraints on $H_{0}$ and $\Omega_{M}$
Figure 3: Typical example of fitting the $D_{L}-z$ relation which is derived from the EMRI dark sirens detectable by the TianQin+LISA network. Here we use the fiducial EMRI population model. The horizontal colored serial squares represent the probability distribution of redshift (normalized by the maximum) derived from the observed candidate host galaxies, and the color hues from white to red show the probability of redshift. The luminosity distances shown here represent the mean values, and their errors ($1\sigma$ CI) are indicated by the black error bars. The solid green line represents the true cosmological model adopted by our simulations. The dot-dashed red line represents the best-fit cosmological model, while the two shaded regions represent the $1\sigma$ and $2\sigma$ CIs.
Figure 4: Constraints on $h$ and $\Omega_{M}$ given by the EMRIs detectable by TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). The horizontal axis represents the different EMRI population models. The error bars correspond to an $1\sigma$ CI. The gray-shaded area represents the $1\sigma$ equivalent error of the prior, and the green-shaded region represents an error scale of $2\%$ ($10\%$) for $h$ ($\Omega_{M}$).
We first fit the $D_{L}-z$ relation, using the luminosity distances from EMRI observations and the redshifts extracted from galaxy surveys. The result is shown in Figure 3, and the corresponding constraints on $H_{0}$ and $\Omega_{M}$ are shown in Figure 4. Note that in this section we fix the parameters of the dark energy EoS to $w_{0}=-1$ and $w_{a}=0$.
For the Hubble-Lemaître constant $H_{0}$, under the fiducial population model, i.e., the M1 model, the expected error is about $8.1\%$ for TianQin, and TianQin I+II can reduce this error to about $4.4\%$. Under the other population models from M2 to M12, the error of $H_{0}$ varies considerably from model to model. The variations are mainly caused by the fact that the detection rates of EMRIs predicted by various models vary considerably (see Babak et al., 2017; Fan et al., 2020). In general, TianQin is able to achieve a precision in the range of about $3.4\%-9.6\%$, while that for TianQin I+II is expected to be in the range of about $1.5\%-7.2\%$.
@@
If TianQin and LISA could form a network, i.e., TianQin+LISA, the constraints on $H_{0}$ and $\Omega_{M}$ are significantly improved in all population models, as is shown by the red error bars in Figure 4. For example, in the M1 model, the errors of $H_{0}$ and $\Omega_{M}$ can shrink to about $1.9\%$ and $20\%$, respectively. Under the other models, the precision of $H_{0}$ can be improved to about $0.4\%-7.1\%$, and the error of $\Omega_{M}$ would reduce to the range of about $13\%-43\%$. For completeness, we summarize the expected precisions on $H_{0}$ and $\Omega_{M}$ for various cases in Table I.
In addition, the importance of improving the sky localization by employing the TianQin+LISA network can be clearly seen in the upper panel of Figure 4. In particular, under Models M3 and M10, the posterior of $h$ provided by a single detector, such as TianQin or TianQin I+II, could be significantly biased with respect to the true value. The bias is related to the paucity of EMRIs that can be precisely localized [see Panel (b) of Figure 2], which increases the fluctuation in the posterior distribution of the redshift of an EMRI. Using TianQin+LISA, however, more EMRIs can be precisely localized, and the precision reduces the fluctuation of the redshift distribution function for each EMRI. We notice that a similar trend can be seen in earlier studies (Laghi et al., 2021; Muttoni et al., 2022; Zhu et al., 2022a, b).